# SpaCE-10: A Comprehensive Benchmark for Multimodal Large Language Models in Compositional Spatial Intelligence

### v1

@@ B.1 More Mulitple-Choice Results
GPT4Scene-Qwen2-VL-7B gpt4scene 24.5 42.9 41.6 32.3 28.0 0.0 0.0 1.5 2.5 26.9 33.8 6.1 0.18
@@
GPT-4o-2024-11-20 gpt4 51.7 63.9 63.2 41.6 37.0 20.3 43.3 40.9 18.8 31.9 51.4 31.0 0.60
Gemini-2.0-flash-thinking gemini 42.2 57.1 40.8 25.5 36.7 41.1 23.2 29.6 6.9 30.6 44.3 26.2 0.59
Claude-3.7-sonnet-20250219 claude3 26.0 44.4 39.3 26.8 45.0 24.3 30.5 31.4 9.4 35.0 36.3 26.1 0.72
@@
InternVL2.5-1B internvl2.5 32.4 22.2 38.0 22.6 28.3 2.0 7.4 7.4 5.9 5.4 28.7 5.6 0.20
InternVL2.5-2B internvl2.5 34.8 29.9 39.8 28.8 17.2 8.9 0.0 8.9 3.5 0.0 30.1 4.2 0.14
Qwen2.5-VL-3B-Instruct qwen2.5 11.0 46.7 50.7 31.6 25.3 13.9 17.2 20.2 11.4 13.1 33.0 15.1 0.46
VILA1.5-3B vila 12.9 31.7 34.6 31.6 35.3 0.0 1.0 2.5 2.0 19.4 29.2 5.1 0.18
InternVL2.5-4B internvl2.5 36.0 50.2 50.7 16.2 51.7 11.9 25.1 25.6 17.3 43.8 40.9 24.7 0.60
Qwen2.5-VL-7B-Instruct qwen2.5 15.0 49.3 37.3 32.3 24.8 17.3 18.7 16.7 17.3 18.8 31.7 17.7 0.56
LLaVA-v1.5-7B liu2023visual 13.3 39.6 44.2 30.1 63.0 0.6 4.9 10.3 6.9 35.0 38.0 11.5 0.30
LLaVA-OneVision-7B llava-onevision 13.1 50.0 50.6 34.5 52.2 4.0 22.7 26.1 10.9 19.4 40.0 16.6 0.41
InternLM-XComposer2.5-7B internlm-xcomposer 6.6 39.4 41.5 34.9 29.8 1.98 22.1 25.6 16.3 20.0 30.4 17.1 0.56
Cambrian-8B cambrian 26.7 34.8 32.6 32.3 41.4 0.5 5.9 12.8 5.9 20.6 33.5 9.1 0.27
VILA1.5-8B vila 15.5 37.0 43.2 24.1 26.7 0.0 0.5 1.5 2.0 26.9 29.3 6.1 0.21
InternVL2.5-8B internvl2.5 14.1 48.1 55.4 33.6 34.5 13.9 23.2 33.5 13.9 11.3 37.3 19.1 0.51
Cambrian-13B cambrian 16.0 27.7 32.8 30.5 31.0 0.5 2.0 10.3 6.4 15.0 27.6 6.8 0.25
CogVLM2-LLaMA3-Chat-19B cogvlm 47.2 43.2 49.3 33.6 64.5 11.8 21.6 22.6 10.3 24.3 47.5 18.1 0.38
Cambrian-34B cambrian 23.4 31.9 27.0 24.7 23.2 0.0 3.9 6.8 6.4 2.5 26.0 3.9 0.15
InternVL2.5-26B internvl2.5 56.7 54.9 54.0 33.0 37.3 29.7 28.1 32.5 17.8 31.3 47.1 27.8 0.59
Qwen2.5-VL-32B-Instruct qwen2.5 11.7 51.3 38.1 32.3 11.4 11.7 23.6 14.8 7.4 42.5 29.1 20.0 0.69
InternVL2.5-38B internvl2.5 58.2 52.7 36.8 31.6 61.9 19.8 46.3 33.5 17.3 43.1 48.2 32.0 0.66
VILA1.5-40B vila 10.1 52.0 58.4 40.9 31.4 11.9 35.5 37.4 19.3 24.3 38.5 25.6 0.67
LLaVA-OneVision-72B llava-onevision 66.6 64.1 62.5 40.3 65.9 34.2 34.5 32.5 16.3 35.0 59.8 30.5 0.51
Qwen2.5-VL-72B-Instruct qwen2.5 26.0 53.7 37.8 32.1 31.1 14.5 27.6 17.7 13.4 30.0 36.1 20.6 0.57
InternVL2.5-78B internvl2.5 54.6 58.8 61.7 40.3 55.8 17.3 35.5 21.3 37.4 50.6 54.2 32.4 0.60

### v2

@@ B.1 More Mulitple-Choice Results
GPT4Scene-Qwen2-VL-7B gpt4scene 38.0 38.9 41.6 29.5 28.0 0.5 1.0 0.0 0.0 19.4 35.2 4.0 0.11
@@
GPT-4o-2024-11-20 gpt4 56.2 58.3 56.2 41.6 52.2 16.8 36.9 41.9 12.7 35.8 52.9 28.8 0.54
Gemini-2.0-flash-thinking gemini 53.1 42.6 53.8 42.2 46.7 47.8 28.6 14.0 6.4 35.6 47.7 26.5 0.56
Claude-3.7-sonnet-20250219 claude3 49.1 46.0 53.8 44.3 49.3 35.0 33.0 13.3 25.7 41.9 48.4 31.3 0.65
@@
InternVL2.5-1B internvl2.5 18.8 43.6 29.9 30.9 41.0 4.4 3.9 4.7 0.5 11.4 32.8 3.0 0.09
InternVL2.5-2B internvl2.5 27.0 36.6 28.8 34.0 48.2 1.3 2.0 4.7 1.5 3.5 34.9 4.1 0.12
Qwen2.5-VL-3B-Instruct qwen2.5 47.1 51.7 31.6 25.5 37.0 21.7 24.1 8.0 15.3 20.9 38.6 17.9 0.46
SpaceOm 47.3 49.7 32.7 21.9 36.7 19.2 24.6 9.3 0.5 20.6 37.7 14.8 0.39
SpaceQwen 41.2 52.3 35.2 28.4 36.4 17.2 26.1 10.0 1.5 35.0 38.7 17.9 0.46
SpaceThinker-Qwen2.5VL-3B 46.7 20.5 33.4 22.4 36.9 22.7 24.1 8.0 1.5 18.8 32.0 15.0 0.47
VILA1.5-3B vila 31.7 34.6 31.6 35.3 12.9 1.0 2.5 2.0 0.0 19.4 29.2 5.0 0.17
InternVL2.5-4B internvl2.5 50.2 50.8 16.2 38.3 56.0 22.2 24.6 4.7 0.5 10.5 42.3 12.5 0.3
Qwen2.5-VL-7B-Instruct qwen2.5 36.9 35.3 32.3 27.6 34.2 20.2 11.8 14.7 13.4 10.4 33.3 14.1 0.42
LLaVA-v1.5-7B liu2023visual 30.5 35.7 22.9 10.7 57.4 4.9 10.3 6.9 0.6 35.0 31.4 11.5 0.37
LLaVA-OneVision-7B llava-onevision 46.4 57.3 34.5 43.3 61.6 21.2 22.7 10.0 5.6 37.8 48.2 19.5 0.40
Cambrian-8B cambrian 34.8 32.6 32.3 25.1 41.4 5.9 12.8 5.9 0.5 20.6 33.2 9.1 0.27
VILA1.5-8B vila 27.5 32.7 17.2 12.4 26.7 0.0 0.5 1.5 2.0 26.9 23.3 6.2 0.27
InternVL2.5-8B internvl2.5 50.0 55.0 33.6 41.1 59.1 8.4 14.8 8.7 12.9 1.5 47.8 10.6 0.2
InternVL2.5-26B internvl2.5 62.6 55.4 33.0 50.2 61.8 30.5 32.0 13.3 17.3 32.8 52.6 25.2 0.49
Qwen2.5-VL-32B-Instruct qwen2.5 48.9 36.8 32.3 31.1 30.1 25.6 6.9 8.0 10.4 12.9 35.8 12.8 0.36
InternVL2.5-38B internvl2.5 64.4 54.3 36.8 45.6 63.0 47.8 37.9 7.3 21.3 46.8 52.8 32.0 0.61
LLaVA-OneVision-72B llava-onevision 67.9 64.5 40.3 46.7 67.3 38.9 32.0 13.3 34.7 35.8 57.3 30.9 0.54
Qwen2.5-VL-72B-Instruct qwen2.5 55.7 40.9 32.1 36.5 38.0 17.2 13.3 13.0 15.7 15.9 40.6 15.0 0.37
InternVL2.5-78B internvl2.5 62.4 64.4 40.3 36.6 67.3 38.4 33.5 12.1 12.9 45.8 54.2 28.5 0.53
@@ B.2 Pipeline Component Ablation
B.3 More Capability Assessment
In this section, we visualize all tested models’ capability performance in Fig. 6, which also presents the same tendency as Fig. 4 that the C5 counting ability is the most difficult for existing models.
Figure 6: Results of all tested MLLMs on 10 atomic capabilities of SpaCE-10. Each value reflects the model’s average accuracy (%) across all question types involving the respective spatial capability (C1–C10), as defined in the benchmark’s task-to-capability mapping.